The marginal product of labour in the production of flash discs is 80 chips per hour. The marginal rate of technical substitution of hours of labour for hours of machine capital is 2.

What is the marginal product of capital?


1
Expert's answer
2020-12-23T07:19:31-0500

The marginal rate of technical substitution is defined as the ratio of marginal product of labor and marginal product of capital.


"MRTS = \frac{MP_{L} }{MP_{K}}"


"2=\frac{80{} }{MP_{K}}"


"MP_{K}=\frac{80}{2} = 40"


The marginal product of capital = 40